A uniform electric field E = 5000 N/C passing through a flat square area A = 2 m2. Determine the electric flux. Theta = 60 degrees ( the angle between the electric field direction and a line drawn perpendicular to the area)

1
Expert's answer
2022-03-01T17:33:33-0500

By definition, the electric flux is:


"\Phi = EA\cos\theta = 5000N/C\cdot 2m^2\cos 60\degree = 5000Nm^2/C"